A box contains 8 balls of the same size but different colours. 4 are red, 3 are blue and 1 is white. If a ball is selected at random,

6.1 What is the probability that a red ball will be selected ? (1)

6.2 What is the probability that a blue or a white ball will be selected? (3)

6.3 What is the probability that a white ball will be selected?


1
Expert's answer
2021-06-30T17:05:31-0400
"4+3+1=8"

6.1


"P(red)=\dfrac{4}{8}=\dfrac{1}{2}"

6.2


"P(blue\ or\ white)=\dfrac{3+1}{8}=\dfrac{1}{2}"



6.3


"P(white)=\dfrac{1}{8}"